OpenStudy (runnergirl17):

Please Help!!!!!!!!! ((cscx/(1+cscx))-((cscx/(1-cscx))

OpenStudy (anonymous):

bunch of algebra that has nothing to do with trig as \(\csc(x)=\frac{\cos(x)}{\sin(x)}\) put \(\cos(x)=a, \sin(x)=b\) and simplify the compound fraction \[\large \frac{\frac{a}{b}}{1+\frac{a}{b}}-\frac{\frac{a}{b}}{1-\frac{a}{b}}\]

OpenStudy (anonymous):

multiply top and bottom by \(b\) and get \[\large \frac{a}{b+a}-\frac{a}{b-a}\]

OpenStudy (anonymous):

then subtract once you have subtracted, go back to sines and cosines

OpenStudy (runnergirl17):

Ok thanks!! I figured it out while I was waiting but I will keep your advice in mind for my other problems :)